# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 4.4.2. Restriction of test vector to components of Shimura set
We describe some consequences of the proof of Theorem 4.14, which will be used in our later arguments.
Let $\varphi\in{\rm M}_{k}(\widehat{R}^{\times},\overline{{\mathbb{Q}}})$ be the test vector as in §4.3.1 associated to $\lambda$.
@@
As in Theorem 4.14, let ${\mathfrak{p}}\nmid 2\ell{\rm{N}}_{K/{\mathbb{Q}}}({\mathrm{Cond}}(\lambda))$ be a degree one prime of $F$ inert in $K$, $\chi_{0}=1$, and ${\mathfrak{c}}=1$. We identify $F_{{\mathfrak{p}}}$ with ${\mathbb{Q}}_{p}$ and the matrices appears in the Hecke action are seen as element of $B_{{\mathfrak{p}}}={\mathrm{GL}}_{2}({\mathbb{Q}}_{p})$. In view of (4.10) and the fact that $T_{\mathfrak{p}}$ has eigenvalue 0 on newforms in the unramified supercuspidal representation $\pi_{\lambda,{\mathfrak{p}}}$, it follows that
$\sum_{t\in(\mathbb{Z}/p^{a}\mathbb{Z})^{\times}}\rho\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\widetilde{H}_{\ell}=p^{a}H_{\ell}+{p^{a-1}}\rho\left(\begin{pmatrix}p^{-2}&\\ &1\end{pmatrix}\right)H_{\ell}.$
@@
$H_{\ell}:=\langle{x,\widehat{\varphi}}\rangle_{x}\pmod{\varpi}$,
where $x=x_{0}$. Therefore,
$p^{-a}\sum_{t\in(\mathbb{Z}/p^{a}\mathbb{Z})^{\times}}\rho\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\widetilde{H}_{\ell}=\langle{x,\left(1+\frac{1}{p}\rho\begin{pmatrix}p^{-2}&\\ &1\\ \end{pmatrix}\right)\widehat{\varphi}}\rangle_{x}\pmod{\varpi}$. (4.11)
Put $\varepsilon$ for the sign of $\varepsilon(\lambda)$. As seen in the proof of Theorem 4.14, $\widetilde{H}_{\ell}^{-\varepsilon}$ is Eisenstein but $\widetilde{H}_{\ell}^{\varepsilon}$ is non-Eisenstein. Moreover, $\sum_{t\in(\mathbb{Z}/p^{a}\mathbb{Z})^{\times}}\rho\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\widetilde{H}_{\ell}$ is non-Eisenstein. Therefore the proof shows that $\sum_{t\in(\mathbb{Z}/p^{a}\mathbb{Z})^{\times}}\rho\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\widetilde{H}_{\ell}^{-\varepsilon}$ is Eisenstein, and in turn $\sum_{t\in(\mathbb{Z}/p^{a}\mathbb{Z})^{\times}}\rho\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\widetilde{H}_{\ell}^{\varepsilon}$ non-Eisenstein.
In view of (4.11) and Lemma 4.10, it follows that $\langle{x,\widehat{\varphi^{\varepsilon}}}\rangle_{x}\pmod{\varpi}$ is non-Eisenstein and $\langle{x,\widehat{\varphi^{-\varepsilon}}}\rangle_{x}$ Eisenstein. Hence $\langle{x,\widehat{\varphi^{-\varepsilon}}}\rangle_{x}\pmod{\varpi}$ is zero by Lemma 4.9.
@@
$\ell\Sigma>k-2\Sigma\in{\mathbb{Z}}[\Sigma]$
so that $\rho_{k,\ell}\pmod{\varpi}$ is an irreducible $G(F_{\ell})$-representation. So the non-vanishing of $\widehat{\varphi^{-\varepsilon}}\pmod{\varpi}$ is equivalent to $\ell$-optimality of $\widehat{\varphi^{-\varepsilon}}\pmod{\varpi}$ with respect to $x$. In view of the preceding paragraph this is a contradiction. ∎
We now describe an application to the split case. Let $\mathfrak{p}$$\nmid\ell$ be a prime of $F$ split in $K$. Let $\chi_{0}$ be a finite order anticyclotomic Hecke character over $K$ such that the condition (D) holds. As in the inert case (4.7), define $\widetilde{H}_{\ell}$ with respect to to $\mathfrak{p}$} and $\chi_{0}$.
@@
Let $\ell$ be a prime such that $(\ell,2D_{K/F})=1$. If $\varepsilon(\lambda)=\pm 1$, then $\widetilde{H}_{\ell}^{\pm}$ is non-Eisenstein and $\widetilde{H}_{\ell}^{\mp}=0$.
@@
By Proposition 4.15, $\langle{x,\widehat{\varphi^{\varepsilon(\lambda)}}}\rangle_{x}\pmod{\varpi}$ is non-Eisenstein and
$\langle{x,\widehat{\varphi^{-\varepsilon(\lambda)}}}\rangle_{x}=0.$
Hence, in view of Lemma 4.7 of and [9, Prop. 3.12] it follows that $H_{\ell}^{\pm}\pmod{\varpi}$ is non-Eisenstein and $H_{\ell}^{\mp}\pmod{\varpi}$ is Eisenstein.
To see the assertion for $\widetilde{H}_{\ell}^{\pm}$, just note that they are related as in the proof of Theorems 4.14 (cf. Equations (4.9) and (4.10)), and then Lemma 4.10 applies. ∎
@@ 4.4.3. Split case
Let $\mathfrak{p}$} be a degree $1$ prime of $F$ split in $K$.
@@
by Lemma 4.11. Our main result is the following.
@@
Let $K$ be a CM field and $F$ the maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\pi_{\lambda}$ the associated cuspidal automorphic representation of conductor ${\mathfrak{N}}_{\lambda}$. Let $\mathfrak{p}$} be a degree one prime of $F$ that is split in $K$ and $\ell$ a prime such that $(\ell,2{\mathfrak{p}}D_{K/F})=1$. Let $\chi_{0}$ be a finite order anticyclotomic Hecke character over $K$ such that the condition (D) holds. Then for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$, we have
${v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)}{\Omega_{\lambda}}\right)=0.$
@@
The argument is similar to the proof of 4.14, whose notation will appear below. In particular, we similarly define $\widetilde{H}_{\ell}$ and $\Theta_{\chi_{0},n}(\pi_{\lambda})$. We may assume that $\varepsilon(\lambda\chi_{0})=+1$ as otherwise $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}<\infty$.
Now under the assumption $\varepsilon(\lambda\chi_{0})=+1$ note that $\widetilde{H}^{\eta_{K/F}(\prod_{v|{\mathfrak{c}}}\varsigma_{v}^{({\mathrm{ord}}_{v}{\mathfrak{c}})})}_{\ell}$ is non-Eisenstein by Corollary 4.16 and the fact that

### v2

@@ 4.4.2. Restriction of test vector to components of Shimura set
We describe some consequences of the proof of Theorem 4.15, which lead to non-Eisenstein-ness of the test vector on specific component of the Shimura set determined by the epsilon factor and will be crucially used in our non-vanishing arguments.
Let $\varphi\in{\rm M}_{k}(\widehat{R}^{\times},\overline{{\mathbb{Q}}})$ be the test vector as in §4.3.1 associated to $\lambda$. Suppose that $\mathfrak{p}$} is coprime to $\mathfrak{N}_{\lambda}^{-}$. Then $\varphi$ is independent of ${\mathfrak{p}}{\mathfrak{c}}$ in the sense of  Remark 3.10. Let $\varphi^{\pm}$ be the restriction of $\varphi$ to $X_{\widehat{R}}^{\pm}$.
@@
As in Theorem 4.15, let $\mathfrak{p}$$\nmid\ell$ be a degree one prime of $F$ inert in $K$, and let $\chi_{0}=1$, and ${\mathfrak{c}}=1$. We also assume $\pi_{\lambda,{\mathfrak{p}}}$ is unramified.
We identify $F_{{\mathfrak{p}}}$ with ${\mathbb{Q}}_{p}$ and the matrices appears in the Hecke action are seen as element of $B^{\times}(F_{\mathfrak{p}})\simeq{\mathrm{GL}}_{2}({\mathbb{Q}}_{p})$ as in §2.2. In view of (4.27) and the fact that $T_{\mathfrak{p}}$ has eigenvalue 0 on newforms in the unramified supercuspidal representation $\pi_{\lambda,{\mathfrak{p}}}$, it follows that
$\sum_{t\in(\mathbb{Z}/p^{s}\mathbb{Z})^{\times}}\rho\left(\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\right)\ H_{\ell,\mu}\equiv p^{s}H_{\ell}+{p^{s-1}}\rho\left(\begin{pmatrix}p^{-2}&\\ &1\end{pmatrix}\right)H_{\ell}\pmod{\varpi}$.
@@
$H_{\ell}:=\langle{x,\widehat{\varphi}}\rangle_{k-2\Sigma},$
where we may and do suppose that $x=x_{0}$ in §3.4 since $\ell$ is auxiliary. Therefore,
$p^{-s}\sum_{t\in(\mathbb{Z}/p^{s}\mathbb{Z})^{\times}}\rho\left(\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\right)H_{\ell,\mu}\equiv\langle{x,\left(1+\frac{1}{p}\rho\left(\begin{pmatrix}p^{-2}&\\ &1\\ \end{pmatrix}\right)\right)\widehat{\varphi}}\rangle_{k-2\Sigma}\pmod{\varpi}$. (4.30)
Put $\varepsilon$ for the sign of $\varepsilon(\lambda)$. As seen in the proof of Theorem 4.15, ${H}_{\ell,\mu}^{-\varepsilon}\pmod{\varpi}$ is Eisenstein but ${H}_{\ell,\mu}^{\varepsilon}\pmod{\varpi}$ is non-Eisenstein. Moreover, $\sum_{t\in(\mathbb{Z}/p^{s}\mathbb{Z})^{\times}}\rho\left(\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\right)H_{\ell,\mu}\pmod{\varpi}$ is non-Eisenstein. Therefore the proof shows that $\sum_{t\in(\mathbb{Z}/p^{s}\mathbb{Z})^{\times}}\rho\left(\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\right)H_{\ell,\mu}^{-\varepsilon}\pmod{\varpi}$ is Eisenstein, and in turn
$\sum_{t\in(\mathbb{Z}/p^{s}\mathbb{Z})^{\times}}\rho\left(\begin{pmatrix}t&\\ &1\\ \end{pmatrix}\right)H_{\ell,\mu}^{\varepsilon}\pmod{\varpi}$
non-Eisenstein.
In view of (4.30) and Lemma 4.11, it follows that $\langle{x,\widehat{\varphi^{\varepsilon}}}\rangle_{k-2\Sigma}\pmod{\varpi}$ is non-Eisenstein and $\langle{x,\widehat{\varphi^{-\varepsilon}}}\rangle_{k-2\Sigma}\pmod{\varpi}$ Eisenstein. Hence $\langle{x,\widehat{\varphi^{-\varepsilon}}}\rangle_{k-2\Sigma}\pmod{\varpi}$ is zero by Lemma 4.10.
@@
$\ell\Sigma>k-2\Sigma\in{\mathbb{Z}}[\Sigma].$
Then it follows from Remark 2.3 that the non-vanishing of $\widehat{\varphi^{-\varepsilon}}\pmod{\varpi}$ is equivalent to $\ell$-optimality of $\widehat{\varphi^{-\varepsilon}}\pmod{\varpi}$ with respect to $x$. In view of the preceding paragraph this is a contradiction. ∎
We now describe an application to the split case. Let $\mathfrak{p}$$\nmid\ell$ be a degree one prime of $F$ split in $K$. Let $\chi_{0}$ be a finite order anticyclotomic Hecke character over $K$ such that the condition (D) holds. As in the inert case (4.24), define ${H}_{\ell,\mu}$ with respect to $\mathfrak{p}$}, $\chi_{0}$, an integer $s\geq 1$ and a primitive character $\mu$ of ${\mathbb{Z}}/p^{s}{\mathbb{Z}}$.
@@
Let $\ell$ be a prime such that $(\ell,2D_{K/F})=1$. If $\varepsilon(\lambda)=\pm 1$, then ${H}_{\ell,\mu}^{\pm}\pmod{\varpi}$ is non-Eisenstein and ${H}_{\ell,\mu}^{\mp}=0$.
@@
Let $\varepsilon\in\{\pm\}$ denote the sign of $\varepsilon(\lambda)$. By Proposition 4.17, $\langle{x,\widehat{\varphi^{\varepsilon}}}\rangle_{k-2\Sigma}\pmod{\varpi}$ is non-Eisenstein and
$\langle{x,\widehat{\varphi^{-\varepsilon}}}\rangle_{k-2\Sigma}=0.$
Since $\mathfrak{c}$} is only divisible by inert primes in the extension $K/F$, in view of (4.17), if $\langle{x,\widehat{\varphi^{-\varepsilon}}}\rangle_{k-2\Sigma}$ vanishes then so does $H_{\ell}^{-\varepsilon}$. Furthermore, in view of Lemma 4.8 and [8, Prop. 3.12], if $\langle{x,\widehat{\varphi^{\varepsilon}}}\rangle_{k-2\Sigma}\pmod{\varpi}$ is non-Eisenstein then so is $H_{\ell}^{\varepsilon}\pmod{\varpi}$.
To see the assertions for ${H}_{\ell,\mu}^{\varepsilon}\pmod{\varpi}$ and ${H}_{\ell,\mu}^{-\varepsilon}$, just note that ${H}_{\ell,\mu}$ and $H_{\ell}$ are related as in the proof of Theorems 4.15 (see equations (4.26) and (4.27)). Since $\mathfrak{p}$} is split in $K$ and $H_{\ell}^{-\varepsilon}=0$, it thus follows that ${H}_{\ell,\mu}^{-\varepsilon}=0$. The assertion for ${H}_{\ell,\mu}^{\varepsilon}\pmod{\varpi}$ then follows by applying Lemma 4.11 to $H_{\ell}^{\varepsilon}\pmod{\varpi}$. ∎
@@ 4.4.3. Split case
Let $\mathfrak{p}$} be a prime of $F$ that is split in $K$.
@@
by Lemma 4.12. It follows that $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}=\infty$ if and only if $\varepsilon(\lambda\chi_{0})=+1$.
Our main result is the following.
@@
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\sigma_{\lambda}$ the associated ${\mathrm{GL}}_{2}(F)$-cuspidal automorphic representation of conductor ${\mathfrak{N}}_{\lambda}$. Let $\mathfrak{p}$} be a degree one prime of $F$ that is split in $K$ and $\ell$ a prime such that $(\ell,2{\mathfrak{p}}D_{K/F})=1$. Let $\chi_{0}$ be a finite order anticyclotomic Hecke character over $K$ such that the condition (D) holds. Then for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$, we have
${v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1/2,\sigma_{\lambda,K}\otimes\chi_{0}\nu)}{\Omega_{\lambda}}\right)=0.$
@@
The argument is similar to the proof of 4.15, whose notation will appear below. In particular, we similarly define ${H}_{\ell,\mu}$ and $\Theta_{\chi_{0},n}(\pi_{\lambda})$. We may assume that $\varepsilon(\lambda\chi_{0})=+1$ as otherwise $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}<\infty$.
Now under the assumption $\varepsilon(\lambda\chi_{0})=+1$ note that ${H}^{\eta_{K/F}(\prod_{v|{\mathfrak{c}}}\varsigma_{v}^{({\mathrm{ord}}_{v}{\mathfrak{c}})})}_{\ell,\mu}\pmod{\varpi}$ is non-Eisenstein by Corollary 4.18 and the fact that